# The Basis of Foot-Sortable Sock Orderings

### v2

@@ 1. Introduction
Consider the following problem, which was introduced by Defant and Kravitz [1]. Starting with an arbitrary line of colored socks to the right of a foot, one aims to line the socks up by color to the left of the foot, but, at any point in time, only one of the following two actions can be performed.
@@
The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if, using only the two specified operations, one can produce a sorted ordering of the socks on the left. Otherwise, it is unsortable. Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering. For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering $rgbrgb$ is sortable (to $rrbbgg$ on the left) while the sock ordering $rgbgrgb$ is unsortable.
Let $X$ and $Y$ be sock orderings. We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly. The subordering $Y\subseteq X$ is proper if $X\neq Y$. An unsortable sock ordering is minimally unsortable if every proper subordering is sortable. We write $a\in X$ if the color $a$ appears in $X$ and we let $X-a$ be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$. We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$. We let $XY$ be the concatenation of the orderings $X$ and $Y$. We say $X$ and $Y$ are equivalent and write $X\cong Y$ if $Y$ can be obtained from $X$ via a bijective mapping of the colors of $X$ to the colors of $Y$. We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$. Otherwise, we say that $X$ avoids $Y$. Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
Modulo the relation $\cong$ and with respect to the partial ordering $\preceq$, Defant and Kravitz called the set of minimal elements among the collection of unsortable sock orderings the basis of foot-sortable sock orderings [1]. One may equivalently view the basis of foot-sortable sock orderings as a collection of pairwise inequivalent minimally unsortable sock orderings that no unsortable sock ordering avoids. In this paper, we give an explicit description of the basis of foot-sortable sock orderings (Section 2.2) and a proof that no unsortable sock ordering avoids this basis (Theorem 3).
@@ 1.1. Prior Work
As observed by Defant and Kravitz [1], this problem fits into the rich line of research dealing with sorting procedures in which the employed data structures are restricted. For example, sorting permutations with a stack was considered by Knuth [3]. Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color. Of particular interest with respect to socks, is the case when there are at most two socks of every color. Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6]. In doing so, Yu determined that this basis is infinite, and this implies that the basis of foot-sortable sock orderings is also infinite. This answered a question of Defant and Kravitz which our work also addresses. Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$. Provided that the ordering is sortable, this algorithm can also procedure a valid sorting. This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
2. Results
2.1. Preliminaries
@@
Throughout, we will refer to the sorting process using primarily stack-centered language. We denote a particular state of the foot-sorting process as $(S,R)$. Here $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot), while $R$ represents what remains on the right. We say that the state $(S,R)$ is foot-sortable (or just sortable) if we can apply the operations stack and unstack so that all of the socks in $SR$ are arranged by color to the left of the foot. Note that if the stack $S$ is not sorted, then $(S,R)$ is unsortable. As such, we implicitly assume throughout that the stack $S$ is sorted.
@@
Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level. At each point, we are mainly concerned with the next sock color that will be moved completely to the left and we say that a color is sortable if it is possible to do this while keeping the stack is sorted. More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
@@
If a color $a\in R$ does not meet the second condition, we say that the color $a$ is blocked by a sandwich. We refer to a sandwich of the form $bcb$ as either a $b$-sandwich or a $c$-sandwich. Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal. Given a sock ordering $X$, a color $a\in X$ is sortable (resp. unsortable) if it is sortable (resp. unsortable) in $(\emptyset,X)$, and a sock ordering $X$ is trivially unsortable if no color in $X$ is sortable. In other words, $X$ is trivially unsortable if the state $(\emptyset,X)$ is terminal.
@@
2.2. The basis of foot-sortable sock orderings
We use $\Gamma$ to denote the basis of foot-sortable sock orderings. As mentioned previously, the set $\Gamma$ is infinite and we use the following definition to describe the infinite classes.
Definition 1.
@@ 2.2. The basis of foot-sortable sock orderings
Observation 2.
Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors. For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable. Furthermore, if $S^{\prime}\subseteq abt_{i}\subseteq S$ and $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$ are such that $S^{\prime}R^{\prime}\neq abt_{i}I(t_{i},\dotsc,t_{n})at_{n}$, then $(S^{\prime},R^{\prime})$ is sortable.
@@
The proof is by induction on $n-i$. For the base case, we have $i=n$, so
@@
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
@@
The following tables describe the basis $\Gamma$. The first table consists of trivially unsortable orderings. It is not difficult to verify that the orderings are minimally unsortable.
@@
Note that all of the sock orderings in the next table have either $atdarat$, $atdrart$, or $atdarbrt$ as a subordering. Furthermore, sorting the color $d$ (preceded or followed by sorting the color $e$ when it appears) for each of these three possible suborderings yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal. Further note that in the orderings below, $d$ and $e$ are the only sortable colors, which implies that the orderings are unsortable. It is then not difficult to verify that the orderings are minimally unsortable.
@@
The next table consists of the sock orderings that have either $abtdabt$ or $abtdatb$ as a subordering. In a similar fashion to the other classes, sorting $d$ (and possibly $e$) first in these orderings yields some $(S,R)$ where $abt\subseteq S$ and $at\subseteq R)$. Since $(abt,at)$ is unsortable and $d$ and $e$ are the only sortable colors in the following orderings, we have that orderings in this class are unsortable, and it is not difficult to verify that these are minimally unsortable.

### v3

@@ 1. Introduction
Consider the following problem, which was introduced by Defant and Kravitz [1]. Starting with an arbitrary line of colored socks to the right of a foot, one aims to move the socks to the left of the foot so that the ordering is sorted by color; that is, the output ordering consists of blocks of socks so that all socks of the same color appear consecutively. At any point in time, however, only one of the following two actions can be performed.
@@
The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if using only the two specified operations, one can produce a sorted ordering of the socks on the left. Otherwise, it is unsortable. Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering. For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering rgbrgbrgbrgb is sortable (to rrbbggrrbbgg on the left) while the sock ordering rgbgrgbrgbgrgb is unsortable.
We will use the following two definitions to describe our result. Two sock orderings are equivalent if one can be obtained from the other by a bijective mapping of the colors. So, for example, we think of rgbgrgbrgbgrgb as equivalent to gbrbgbrgbrbgbr. We call a sock ordering minimally unsortable if it is unsortable and removing any sock yields a sortable sock ordering. For example, rgbgrgbrgbgrgb is minimally unsortable, because it is unsortable and the sock orderings gbgrgbgbgrgb, rbgrgbrbgrgb, rggrgbrggrgb, rgbrgbrgbrgb, rgbggbrgbggb, rgbgrbrgbgrb, and rgbgrgrgbgrg are each sortable (to gggrbbgggrbb, rrggbbrrggbb, rrgggbrrgggb, rrbbggrrbbgg, rgggbbrgggbb, ggbbrrggbbrr, and bgggrrbgggrr, respectively).
Modulo equivalence, Defant and Kravitz called the unique set of minimally unsortable sock orderings the basis of foot-sortable sock orderings [1]. In this paper, we explicitly described this set. This description appears in Section 2.3, and the rest of the paper is devoted to proving that this set is indeed the basis of foot-sortable sock orderings. (For reference, rgbgrbgrgbgrbg is equivalent to the ordering $\mathcal{T}_{6}$ that is defined in Section 2.3.)
@@ 1.1. Prior Work
As observed by Defant and Kravitz [1], this problem fits into the rich line of research dealing with sorting procedures in which the employed data structures are restricted. For example, sorting permutations with a stack was considered by Knuth [3]. Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color. Of particular interest with respect to socks is the case when there are at most two socks of every color. Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6]. In doing so, Yu determined that this basis is infinite, and this implies that the basis of foot-sortable sock orderings is also infinite. This answered a question of Defant and Kravitz which our work also addresses. Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$. Provided that the ordering is sortable, this algorithm can also produce a valid sorting. This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
2. Preliminary discussion and a description of the basis.
2.1. Additional Definitions and Notation
Let $X$ and $Y$ be sock orderings. We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly. The subordering $Y\subseteq X$ is proper if $X\neq Y$. So, an unsortable sock ordering is minimally unsortable if every proper subordering is sortable. We write $a\in X$ if the color $a$ appears in $X$ and we let XaX-a be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$. We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$. We let $XY$ be the concatenation of the orderings $X$ and $Y$ and we write $X\cong Y$ when $X$ and $Y$ are equivalent. We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$. Otherwise, we say that $X$ avoids $Y$. Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
@@ 2.1. Additional Definitions and Notation
Throughout, we will refer to the sorting process using primarily stack-centered language. We denote a particular state of the foot-sorting process as $(S,R)$. Here $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot), while $R$ represents what remains on the right. We say that the state $(S,R)$ is foot-sortable (or just sortable) if we can apply the operations stack and unstack so that all of the socks in $SR$ are arranged by color to the left of the foot. Note that if the stack $S$ is not sorted, then $(S,R)$ is unsortable. As such, we implicitly assume throughout that the stack $S$ is sorted. We call a state $(S^{\prime},R^{\prime})$ a substate of $(S,R)$ if $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$, and it is proper if at least one of the two containment relations is strict.
@@
Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level. At each point, we are mainly concerned with the next sock color that will be moved completely to the left, and we say that a color is sortable if it is possible to do this while keeping the stack sorted. Of great importance in the context of moving a color to the left is the concept of a sandwich: a configuration of (possibly non-consecutive) socks taking the form bcbbcb, where $b$ and $c$ are distinct colors. We refer to a sandwich of the form bcbbcb as either a $b$-sandwich or a $c$-sandwich. More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
@@
If a color $a\in R$ does not meet the second condition for colors $b,c$ in $SR$, we say that the color $a$ is blocked by the sandwich bcbbcb. Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal. Given a sock ordering $X$, a color $a\in X$ is sortable (resp. unsortable) if it is sortable (resp. unsortable) in $(\emptyset,X)$, and a sock ordering $X$ is trivially unsortable if no color in $X$ is sortable. In other words, $X$ is trivially unsortable if the state $(\emptyset,X)$ is terminal.
@@
2.2. Interlaced sequences
As mentioned previously, the basis of foot-sortable sock orderings is infinite. We will use the following definition to help describe the infinite classes.
@@ 2.2. Interlaced sequences
Observation 1.
Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors. For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable. Furthermore, for any $0\leq i\leq n$, if $(S^{\prime},R^{\prime})$ is a proper substate of $(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})$, then $(S^{\prime},R^{\prime})$ is sortable.
@@
The proof is by induction on nin-i. For the base case, we have $i=n$, so
@@
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that when ni=1n-i=1 the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
@@
2.3. The basis of foot-sortable sock orderings
The tables below describe the basis of foot-sortable sock orderings. We use $\Gamma$ to denote this set, so our main theorem is the following.
Theorem 2.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
With Theorem 2, to show that $\Gamma$ is the basis of foot-sortable sock orderings we also need to show that every ordering in $\Gamma$ is minimally unsortable.
The following table consists of the trivially unsortable orderings in $\Gamma$.
@@ 2.3. The basis of foot-sortable sock orderings
It is not hard to verify that for each ordering listed above, every color is blocked by sandwich, so the orderings are indeed trivially unsortable. It is also not difficult to verify that whenever a sock is removed, a color becomes sortable. This is enough to show that $\mathcal{T}_{1},\dotsc,\mathcal{T}_{7}$ are minimally unsortable, because they each contain only three distinct colors, which means that after a color is sorted in one of the suborderings only at most two colors remain, but three distinct colors are necessary for a color to be blocked by a sandwich. Similarly, for $\mathcal{T}_{8}$, $\mathcal{T}_{9}$, or $\mathcal{T}_{10}$, it can be checked that removing one sock produces an ordering in which two colors can be successively sorted. After this, only two distinct colors then remain, so the suborderings are sortable.
All of the sock orderings in the next table have either atdaratatdarat, atdrartatdrart, or atdarbrtatdarbrt as a subordering. Each of these three suborderings is sortable if one first sorts the color $a$. However, for these three suborderings, sorting the color $d$ yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal. So, to construct the unsortable sock orderings in this table, either a new sock is inserted so that a $d$-sandwich blocks $a$, or $ed$ is placed immediately after $d$, After this addition, the only sortable colors are $d$ and if $ed$ was added, $e$. This implies that the orderings are unsortable.
@@
To see that these orderings are minimally unsortable, first consider the suborderings of $\mathcal{C}_{1,3}\cong atdatrat$ formed by removing a single sock:
$tdatrat,adatrat,atatrat,atdtrat,atdarat,atdatat,atdatrt,\text{ and }atdatra.$
For adatratadatrat, atatratatatrat, and atdaratatdarat, we have removed a sock from the $d$-sandwich that blocks $a$, so we can fully sort each of these three subordering by first sorting $a$, and then sorting $r$. For the remaining suborderings (tdatrattdatrat, atdtratatdtrat, atdatatatdatat, atdatrtatdatrt, and atdatraatdatra), we can first sort $d$ and then one additional color ($a$, $r$, $t$, $t$, and $t$, respectively), leaving at most two colors, so the suborderings are sortable. For every other subordering formed by removing one sock from an ordering in the table, one can verify that one of the following two possibilities hold:
  • One of the socks in the $d$-sandwich that blocks $a$ has been removed and the subordering can then be fully sorted when one first sorts the color $a$, or
  • sorting $d$ (and then $e$ if it appears) yields either $(S,R)$ or $(S,tR)$ where $(S,R)$ is a proper substate of one of the states $(at,arat)$, $(at,rart)$, or $(at,arbrt)$, thereby implying that the subordering is sortable.
The next table consists of the sock orderings that have either abtdabtabtdabt or abtdatbabtdatb as a subordering.